OpenStudy (anonymous):

HELP a boat traveled 210 miles downstream and back. The trip downstream took 10 hours. The trip back took 70 hours. What is the speed of the boat in still water? What is the speed of the current? i need to graph it please help

OpenStudy (anonymous):

set up a system of equations each trip down stream and upstream is 105 miles each since 210/2 = 105 the trip downstream took 10 hours, so the boat was traveling 105/10 = 10.5 mph downstream the trip upstream took 70 hours, so the boat was traveling 105/70 = 1.5 mph upstream set up a system of equations to find the speed of the boat in still water we'll let x = the speed of the boat in stillwater and let y = the speed of the current downstream x + y = 10.5 (for downstream) x - y = 1.5 (for upstream) solve the system of equations. you can also solve by drawing a graph with these two equations

OpenStudy (anonymous):

im so confused.. my x is the speed of the boat and my y is the speed of the current